4 moles of helium are mixed with three moles of hydrogen to form a mixture molar mass of helium is 4g formed and that of hydrogen is 2 gram equivalent degree of freedom of the mixture is

Grade:11

1 Answers

Arun
5 years ago
 
Cv = 3*5R/2 + 4*3R/2    =27R/2
Cp = Cv + R = 27R/2 + R  =29R/2
now you can calculate equivalent degree of freedom